OpenStudy (anonymous):

What do you get for.. (x - 9)^2 = 1??

myininaya (myininaya):

How do you undo the square?

myininaya (myininaya):

What action must you perform to both sides in order to undo that square on the left?

OpenStudy (anonymous):

Square root

myininaya (myininaya):

Yep. You need plus or minus.

OpenStudy (anonymous):

You must solve it. (x-9)^2 = 1 √Both sides.. √(x-9)^2 = √1 x-9 = √1 since √1 is really just 1... x - 9 = 1 +9 +9 ------------ x = 10

OpenStudy (anonymous):

A plus or minus, for what?

myininaya (myininaya):

So you have \[x-9=\pm 1 \text{ correct?}\]

OpenStudy (anonymous):

Yes x = 10,8

myininaya (myininaya):

@chandhooguy Please in the future let the asker come to the answer themselves. Let them do some work. Thanks.

OpenStudy (anonymous):

How would it equal '10,8'?

myininaya (myininaya):

You have two equations x-9=1 or x-9=-1 Do you know how to solve either of these?

OpenStudy (anonymous):

How do you get two equations..?

myininaya (myininaya):

Because we have absolute value

myininaya (myininaya):

\[\sqrt{x^2}=|x|\]

OpenStudy (anonymous):

What's that supposed to be

myininaya (myininaya):

=x when x>=0 =-x when x<0